if a sector of a circle with an angle of 60° has an area of 24 pi what is the radius of the circle
Answer by jorel555(1290)

60/360=1/6
1/6 Area of circle=24 pi
A=24/(1/6)=144 pi;
pi*r^2=144 pi
r^2=144
r=12
The radius is 12!!!!!!!!!!!!!!!!!